OpenStudy (anonymous):

Determine whether f : Z x Z --> Z is onto if f (m, n) = m + n + 1. Two parts: first, prove or disprove the function is one-to-one AND second, prove or disprove the function is onto.

OpenStudy (anonymous):

Hi Haichi

myininaya (myininaya):

Injective f(a,b)=f(c,b) see if you have a=c and b=d If you do then it is injective.

OpenStudy (anonymous):

Hello, sorry I was away from the computer

OpenStudy (anonymous):

how do I prove if it is one to one function?

myininaya (myininaya):

basically we want to suppose f(a,b)=f(c,d) and if we show (a,b)=(c,d) then we have shown it is injection . Or we can try to think of an example that shows it isn't injective. Can you think of (a,b) not equal to (c,d) such that f(a,b) equal f(c,d)

OpenStudy (anonymous):

f (m, n) = m + n + 1? how do I use (a,b) = (c,d) equation?

myininaya (myininaya):

f(a,b)=f(c,d) a+b+1=c+d+1 a+b=c+d Does this imply a=c and b=d? Well it could actually imply the other way around couldn't it? a=d and b=c So maybe we can try to think of a counterexample: think of choosing one where a=d and b=c

myininaya (myininaya):

(a,b) is what we are plugging in (c,d) is what we are plugging in

OpenStudy (anonymous):

oh ok so any integer will match?

myininaya (myininaya):

I'm saying you can provide a counterexample to the injective part

myininaya (myininaya):

And I gave you a hint as to what to pick

OpenStudy (anonymous):

that will only work if I pick one integer for all a,b,c,d?

myininaya (myininaya):

I will copy what I pasted above: "So maybe we can try to think of a counterexample: think of choosing one where a=d and b=c"

OpenStudy (anonymous):

I am sorry. I dont know what counterexample to use for this question

myininaya (myininaya):

I'm asking you to think of one where a=d and b=c

OpenStudy (anonymous):

to prove that it is incorrect?

myininaya (myininaya):

We are proving this it is not injective.

myininaya (myininaya):

Because we can show that (a,b) not equal to (c,d) implies f(a,b)=f(c,d) for some values a,b,c, and d.

myininaya (myininaya):

so if i want a=d and b=c I will randomly select two points such that that happens (4,3) and (3,4)

myininaya (myininaya):

now what is f(4,3) and what is f(3,4)?

OpenStudy (anonymous):

right so it is not injective?

OpenStudy (anonymous):

im sorry im not sure what you're asking

myininaya (myininaya):

remember that f(m,n)=m+n+1?

OpenStudy (anonymous):

ok so that is not one to one function?

myininaya (myininaya):

Right because we were about to find (a,b) not equal (c,d) such that f(a,b)=f(c,d)

myininaya (myininaya):

(3,4) doesn't equal (4,3) but f(3,4)=f(4,3)=8

myininaya (myininaya):

That is called a counterexample

OpenStudy (anonymous):

ok what about onto function?

myininaya (myininaya):

You want to suppose b is an element of the codomain which happens to be the set of integers in this case. You want to prove there exists (m,n) in the ZxZ for which f(m,n)=b

myininaya (myininaya):

f(m,n)=m+n+1=b We want to chose an m and n such that does =b

myininaya (myininaya):

I will let you choose m and n such that when you simplify m+n+1 you get b

OpenStudy (anonymous):

b would be in co domain? does it have to match domain numbers?

myininaya (myininaya):

what does that mean exactly?

OpenStudy (anonymous):

if (m,n) is in domain and m + n +1 is in codomain, these wont match?

OpenStudy (anonymous):

sorry im just confused lol

myininaya (myininaya):

I kinda don't understand what you say saying exactly I'm just asking you to choose an m and n such that when simplify m+n+1 you get b

OpenStudy (anonymous):

if I chose (3,2) and b would be 6?

myininaya (myininaya):

well the problem with that is you aren't showing you get all the integers

OpenStudy (anonymous):

that would be 5 = 6 in form of (m,n) = m + n + 1?

myininaya (myininaya):

m+n+1=b solving for either m or n m=b-n-1 So I will choose m=b-n-1 and n=n so f(b-n-1,n)=(b-n-1)+n+1=b so this shows that every element of the codomain does get hit Therefore the function is surjective

OpenStudy (anonymous):

f(b-n-1,n)=(b-n-1)+n+1=b ?

myininaya (myininaya):

-n+n=0 -1+1=0 only thing left is b

OpenStudy (anonymous):

how do you know the difference between injective and surjective?

OpenStudy (anonymous):

I understand them by the diagrams but not in the equations

myininaya (myininaya):

f:A->B injective: For every x,y in A..if x does not equal y then f(x) does not equal f(y) --as you seen we found a counterexample for this above. we found x does not equal y such that we had f(x) equals f(y) surjective: If for every b in B there is a in A with f(a)=b ---This is an exisential proof. All we have to is find an element a in the domain such that we get f(a)=b where b represents every element of the codomain B. m and n were both integers b was an integer choosing m to b-n-1 was fine because b-n-1 is also an integer for any integer b b is an integer the set of integers is closed under addition/subtraction

OpenStudy (anonymous):

Ok, i think I get it. Thank you for your help!

myininaya (myininaya):

So in others words for this problem that wasn't the only way to choose m and n.

myininaya (myininaya):

we could have chose to left m as is and to select n as b-m-1

myininaya (myininaya):

of we could have even chose m as 0 and n as b-1

myininaya (myininaya):

or chose m as b-1 and n as 0

myininaya (myininaya):

We just needed to show we get every element of the codomain

myininaya (myininaya):

for the surjective part

OpenStudy (anonymous):

Ok I think I understand it now. Thank you!